OpenStudy (anonymous):

Please help. Does the rule y = 4 * 4^x represent a linear or an exponential function? Explain.

jimthompson5910 (jim_thompson5910):

When you graph out y = 4*4^x, you'll see that a curve results and NOT a straight line. So y = 4*4^x is NOT a linear function.

jimthompson5910 (jim_thompson5910):

So by the process of elimination, this is an exponential function. It turns out that y = 4 * 4^x is an exponential function since it's in the form y = a*b^x where a and b are constants

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome